# Less LLM, More Documents: Searching for Improved RAG

### v2

@@ Abstract
Retrieval-Augmented Generation (RAG) couples document retrieval with large language models (LLMs). While scaling generators improves accuracy, it also raises cost and limits deployability. We explore an orthogonal axis: enlarging the retriever’s corpus to reduce reliance on large LLMs. Experimental results show that corpus scaling consistently strengthens RAG and can often serve as a substitute for increasing model size, though with diminishing returns at larger scales. Small- and mid-sized generators paired with larger corpora often rival much larger models with smaller corpora; mid-sized models tend to gain the most, while tiny and large models benefit less. Our analysis shows that improvements arise primarily from increased coverage of answer-bearing passages, while utilization efficiency remains largely unchanged. These findings establish a principled corpus–generator trade-off: investing in larger corpora offers an effective path to stronger RAG, often comparable to enlarging the LLM itself.
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) [6, 17] combines document retrieval with large language models (LLMs), and has become a popular paradigm for open-domain question answering (QA) [8, 28, 30]. Most prior work has focused on scaling up the generator, which indeed improves accuracy but requires very large and expensive LLMs [7, 19, 21].
In contrast, the retriever controls the external evidence supplied to the generator, thereby directly influencing factuality and mitigating hallucinations [17, 25]. However, the relationship between retriever capacity and generator size remains underexplored. In particular, it is not well understood whether enlarging the retrieval corpus can reduce reliance on larger generators, which is an important question for practical deployment, where smaller LLMs are easier to serve [31, 34, 37].
To address this gap, we conduct a systematic study of the trade-off between corpus scale and generator size by combining randomly-shardable retrieval over ClueWeb22 [20] with open-source Qwen3 models [35] of different scales.
Our experiments on multiple QA benchmarks [1, 13, 16] reveal that enlarging the retrieval corpus not only improves the performance of smaller LLMs, but also enables them to rival or even surpass larger counterparts. For example, a 1.7B-parameter model with a $4\times$ larger corpus outperforms a 4B model, and a 4B model with only a $2\times$ larger corpus consistently outperforms an 8B model.
These findings highlight a practical trade-off: scaling the retrieval corpus can partially substitute for scaling the generator. This insight suggests an efficient and deployable RAG design, where expanding the corpus offers a promising alternative to enlarging the LLM itself.
@@ 2 Related Work
In parallel, retrieval-augmented models have emerged as an alternative path to scaling. Retrieval-augmented language models (RALMs) such as RETRO [2] and Atlas [12] demonstrate that enlarging inference-time datastores consistently improves performance: relatively small generators paired with massive retrieval memory can outperform much larger LM-only baselines. Shao et al. [24] further confirmed this monotonic trend. Unlike modular RAG, which decouples retriever and generator, RALMs integrate retrieval through pretraining-time vector memories, requiring retriever-generator co-training.
@@ 3 Methodology
To address this gap, we present a systematic framework for analyzing corpus–generator trade-offs in RAG. Our approach evaluates whether, and under what conditions, scaling the retrieval corpus can compensate for smaller LLMs. We organize the methodology around two complementary dimensions. First, we study corpus scaling as compensation: does enlarging the retriever corpus enable smaller generators to rival or surpass larger models by leveraging broader retrieval evidence? Second, we investigate differential effects across LLMs: how do models of different sizes benefit from corpus expansion, and are there consistent patterns in how scaling interacts with model capacity? These two dimensions form the backbone of our experimental design and subsequent analysis.
@@ 3.1 Retriever: Corpus Scaling
Let $\mathcal{C}$ denote a fixed corpus. We simulate corpus scaling via a balanced random partition of $\mathcal{C}$ into $N$ disjoint shards $\{S_{1},\ldots,S_{N}\}$ of approximately equal size:
$\Pi(\mathcal{C})\;\to\;\{S_{1},S_{2},\ldots,S_{N}\},\quad S_{i}\cap S_{j}=\varnothing\;\;\forall\,i\neq j,\quad\bigcup_{i=1}^{N}S_{i}=\mathcal{C}$
A corpus scale $n\in\{1,\ldots,N\}$ is realized by activating $n$ shards; without loss of generality we use the canonical prefix $\mathcal{C}^{(n)}=\bigcup_{i=1}^{n}S_{i}$, so that $|\mathcal{C}^{(n)}|\propto n$. For a query $q$, the retriever operates on $\mathcal{C}^{(n)}$ to retrieve top-$k$ documents, transform them into chunks, rerank, and pass the top $m$ chunks to the generator. All retrieval hyper-parameters and downstream settings are fixed across $n$.
@@ 3.2 Generator
We consider a family of generators $\{M_{x}\}$, where $M_{x}$ denotes a model with parameter size $x$ drawn from the same architecture. Each generator takes as input a fixed template consisting of the query and retrieved chunks. Prompting and decoding configurations are kept constant across all $M_{x}$, ensuring that model size is the sole varying factor.
@@ 3.3 Trade-off Formalization
We adopt a full-factorial design pairing each corpus scale $n\in\{1,\ldots,N\}$ with each generator $M_{x}$. For every $(n,x)$ pair, retrieval and decoding settings are fixed, so that only the corpus size $n$ and the model size $x$ vary. Let $P_{m}(n,x)$ denote the evaluation score under metric $m\in\{\text{F1},\text{EM}\}$.
@@
$n^{\star}(x_{small}\!\to\!x_{large})\;:=\;\min_{m\in\{\text{F1},\,\text{EM}\}}\;\min\big\{n\,:\,P_{m}(n,x_{small})\,\geq\,P_{m}(1,x_{large})\big\}$
the smallest corpus scale where a smaller generator $M_{x_{small}}$ matches the 1-shard baseline ($n=1$) of a larger generator $M_{x_{large}}$. We report $n^{\star}$ and efficiency curves across $(n,x)$, jointly characterizing corpus–generator trade-offs without relying on model-specific tricks. Section 4 details datasets, metrics, and constants.
@@ 4.2 Evaluation Metrics
We report F1 and Exact Match (EM) scores, following official evaluation scripts with gold answers. Our analysis primarily focuses on these two metrics throughout the paper.
@@ 4.3 Retriever: Implementation Details
Encoder and index. We use MiniCPM-Embedding-Light [10] for dense passage encoding, selected for its balance between retrieval quality and computational efficiency at web scale. Indexing is performed using DiskANN [26], a widely adopted ANN backend that supports fast multi-shard retrieval, where a similar configuration has also been used in other large-scale retrievers built on ClueWeb22-A [5].
Retrieval pipeline. For each corpus scale $n$, the retriever operates over the active shards to select the top-10 documents, which are segmented into overlapping chunks and reranked. From these, the top-8 chunks are passed to the generator. A higher-capacity reranker from the same embedding family is used, and all retrieval and reranking settings are held fixed across settings to isolate corpus scaling effects.
@@ 4.4 Generator: Implementation Details
We instantiate $\{M_{x}\}$ using the open-source Qwen3 family [35]: Qwen3-0.6B, 1.7B, 4B, 8B, and 14B. This series spans over an order of magnitude in parameter scale, enabling a controlled study of size-related trends.
All models share identical prompting templates and decoding settings. This setup fixes the retrieval pipeline and input assembly while varying only generator scale, allowing a clean measurement of corpus–model trade-offs. While our preliminary experiments with Qwen2.5 and early LLaMA models yielded robustly consistent conclusions, we did not adopt them due to the lack of a homogeneous, wide-ranging family comparable to Qwen3 and slower inference speed.
@@ 5.1 The Effect of Corpus Scaling as Compensation
We ask whether enlarging the retriever corpus can compensate for smaller LLMs, allowing them to match or surpass larger generators in RAG performance. Concretely, we define the retriever’s corpus size $|C|$ in terms of the number of active shards $n$, where each shard indexes $\sim$22M documents from ClueWeb22-A. Thus, corpus scaling corresponds to increasing $n$, and we study how varying $n$ interacts with different LLM sizes.
To investigate this question, we fixed the generator to one of five Qwen3 variants ($M_{0.6B}$, $M_{1.7B}$, $M_{4B}$, $M_{8B}$, $M_{14B}$) and varied corpus scale $n$ by cumulatively activating retrieval shards. For each model, we evaluated the same set of $n$ under a uniform protocol, enabling consistent cross-model comparison.
@@
$\times$7 34.29 23.46 41.50 30.53 46.84 34.20 48.49 35.90 49.39 36.91
$\times$8 35.16 24.56 41.87 31.18 46.89 34.01 48.11 35.88 49.60 37.18
$\times$9 34.82 23.91 41.79 30.70 46.77 33.80 48.51 35.84 49.71 37.20
$\times$10 35.59 24.94 41.82 30.37 46.88 34.22 48.47 35.75 50.18 37.67
$\times$11 35.44 24.97 42.28 30.86 46.91 34.03 48.99 35.90 50.39 37.54
$\times$12 35.96 25.33 42.41 30.92 47.30 34.20 48.35 35.10 50.47 37.75
@@
Our results show clear evidence that corpus expansion enables smaller models to match or even outperform larger counterparts. On NQ, as shown in Table 1, we find that the smallest model needs more corpus to surpass the larger model $n^{\star}(0.6\text{B}\!\to\!1.7\text{B})=5$. For larger generators, compensation is much easier: $n^{\star}(4\text{B}\!\to\!8\text{B})=2$ and $n^{\star}(8\text{B}\!\to\!14\text{B})=2$. These results indicate that, under our setup, scaling corpus size can be a more effective and efficient lever than simply scaling LLM size. Figures 2 and 2 visualize these catch-up points.
@@
$n^{\star}$ NQ TriviaQA WebQ
@@
The same trend holds on TriviaQA and WebQ (Table 2). In the tiny-model regime, corpus scaling is inefficient: e.g., $n^{\star}(0.6\text{B}\!\to\!1.7\text{B})=10$ and $n^{\star}(1.7\text{B}\!\to\!4\text{B})=7$ on TriviaQA. This indicates that corpus scaling is less efficient in the tiny-model regime. In contrast, once the generator reaches medium to large scale, only doubling the corpus is typically sufficient to catch up with the next-tier model. For instance, we find that $n^{\star}(4\text{B}\to 8\text{B})=2$ and $n^{\star}(8\text{B}\to 14\text{B})=2$ on NQ and TriviaQA, and at most $n^{\star}=3$ on WebQ. Detailed results for TriviaQA and WebQ are provided in Table 3 and  4 within the Appendix.
@@
Shards in our corpus are balanced in size but not perfectly uniform, because randomization was applied locally during assignment rather than globally. We also observed early performance saturation in the first few increments of corpus scaling. To probe the sensitivity of generation performance to corpus quality, we reversed the shard order at retrieval time, replacing $S_{1},\ldots,S_{n}$} with $S_{N-n+1},\ldots,S_{N}$}. As expected, this led to a modest decrease in absolute RAG scores (Figure 3, left).
@@
In other words, the lower average corpus quality shifts the performance level downward, yet the relative additional corpus needed for a smaller model to catch up with the next larger model remains essentially stable. This reinforces our earlier conclusion: While corpus quality affects absolute accuracy, scaling corpus quantity still enables smaller generators to overtake larger ones with similar amounts of additional context. Based on this observation, we report all subsequent results using the forward corpus order for consistency.
@@ 5.1.3 Why Does Corpus Scaling Improve RAG?
At the micro level, corpus scaling increases the likelihood that retrieved passages explicitly contain the gold answer. With a small corpus scale ($n=1$), retrieved chunks often lack factual mentions. The larger $n$ brings both the query and the answer terms into the context, directly providing the generator with grounding evidence as intended in RAG.
@@
Summary of Findings.
Across benchmarks, corpus scaling consistently enables smaller generators to catch up with larger ones, with $n^{\star}$ thresholds stable even under lower-quality (reversed) corpora. The mechanism is straightforward: enlarging the corpus raises the probability that retrieved chunks contain gold answers, thereby providing models with comparable evidence. Thus, corpus expansion is a reliable lever for improving RAG effectiveness.
@@ 5.2.1 Classification Methodology
which captures the additional fraction of initially unanswerable questions newly resolved at shard $n$.
@@
Our analysis, summarized in Figure 6, reveals a consistent and model–invariant pattern for how the retrieval scale helps answer questions that are initially unanswerable without context, with similar trends observed on TriviaQA and WebQ (see the Appendix).
@@
5.2.2 Initial Jump, Subsequent Growth, and Saturation
@@
Across all LLM sizes, corpus scaling yields a qualitatively similar CB pattern: a sharp first jump, sustained gains up to roughly $n\!\approx\!6$, and diminishing returns thereafter. The per-shard increments $\Delta_{n}$ also follow a nearly identical pattern across models: peaking early and tapering to near zero (Figure 7). These shared patterns suggest a size-invariant retrieval effect: additional shards do not yield systematically greater CB gains for larger generators than for smaller ones. In practice, corpus expansion mainly shifts CB upward without altering its growth curve, reinforcing our conclusion that scaling the corpus is a robust and efficient lever for RAG.
@@
As shown in Figure 8, the Utilization Ratio remains approximately constant across corpus scales and is clustered across models within a narrow band, indicating stability with respect to $n$. Although both $\mathrm{CB}@n$ and $\mathrm{Coverage}@n$ grow with $n$, their ratio fluctuates only slightly. Thus, corpus scaling primarily raises the coverage of relevant evidence, while the efficiency with which generators convert available evidence into correct answers remains largely unchanged. Consequently, the benefits realized by the generator scale approximately in proportion to the availability of the answer-bearing context, making corpus growth a reliable axis of improvement in RAG.
@@
Summary of Findings.
Across model sizes, corpus scaling follows a common profile: a sharp initial rise followed by gradual gains toward saturation. Further analysis of Utilization Ratio shows that the improvements in RAG stem mainly from increased gold answer coverage, rather than differences in context utilization efficiency between models. This indicates that corpus expansion is a reliable and size-agnostic lever to improve performance.

### v3

@@ Abstract
Retrieval-Augmented Generation (RAG) couples document retrieval with large language models (LLMs). While scaling generators often improves accuracy, it also increases inference and deployment overhead. We study an orthogonal axis: enlarging the retriever’s corpus, and how it trades off with generator scale. Across multiple open-domain QA benchmarks, corpus scaling consistently strengthens RAG and can in many cases match the gains of moving to a larger model tier, though with diminishing returns at larger scales. Small- and mid-sized generators paired with larger corpora often rival much larger models with smaller corpora; mid-sized models tend to gain the most, while tiny and very large models benefit less. Our analysis suggests that these improvements arise primarily from increased coverage of answer-bearing passages, while utilization efficiency remains largely unchanged. Overall, our results characterize a corpus–generator trade-off in RAG and provide empirical guidance on how corpus scale and model capacity interact in this setting.
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) [6, 17] augments large language models (LLMs) with retrieved evidence and is widely used for open-domain question answering (QA) [8, 28, 30]. While much prior work improves RAG by scaling the generator, larger models typically incur higher inference and deployment overhead [7, 19, 21]. In parallel, the retriever determines what evidence is available to the generator and can help mitigate hallucinations [17, 25], yet how retriever-side scaling interacts with generator scale is not well understood.
We study this interaction by varying the retrieval corpus scale while holding the retrieval pipeline and evidence budget fixed. Concretely, we combine randomly-shardable retrieval over ClueWeb22 [20] with open-source Qwen3 models [33] spanning 0.6B–14B parameters, and evaluate on NQ, TriviaQA, and WebQ [1, 13, 16]. Our results show that enlarging the corpus consistently improves RAG and can, in this setting, partially offset smaller generators: for example, a 1.7B model with a $4\times$ larger corpus outperforms a 4B model, and a 4B model with a $2\times$ larger corpus consistently outperforms an 8B model.
Overall, our work characterizes a corpus–generator trade-off in RAG and provides empirical guidance on (i) when corpus scaling can match gains from scaling the generator, (ii) which model-size regimes benefit most from additional corpus, and (iii) why these gains arise, via analyses of answer-bearing evidence coverage and context utilization.
@@ 2 Related Work
In parallel, retrieval-augmented models have emerged as an alternative path to scaling. Retrieval-augmented language models (RALMs) such as RETRO [2] and Atlas [12] demonstrate that enlarging inference-time datastores consistently improves performance: relatively small generators paired with massive retrieval memory can outperform much larger LM-only baselines. Shao et al. [24] further confirmed this monotonic trend. Unlike modular RAG, which decouples retriever and generator, RALMs integrate retrieval through pretraining-time vector memories, requiring retriever-generator co-training. In contrast, we focus on modular RAG with a fixed evidence budget and study corpus–generator trade-offs across model-size regimes, rather than monotonic retrieval scaling in integrated RALMs.
@@ 3 Methodology
We present a systematic framework to analyze corpus–generator trade-offs in RAG, asking when scaling the retrieval corpus can compensate for smaller LLMs and how this interaction varies across model sizes. Our design centers on two questions: (i) corpus scaling as compensation—whether enlarging the corpus enables smaller generators to match or surpass larger ones; and (ii) differential effects across LLMs—how benefits from corpus expansion change with model capacity.
@@ 3.1 Retriever: Corpus Scaling
Let $\mathcal{C}$ denote a fixed corpus. We simulate corpus scaling by randomly partitioning $\mathcal{C}$ into $N$ disjoint shards $\{S_{1},\ldots,S_{N}\}$ of approximately equal size:
$\Pi(\mathcal{C})\;\to\;\{S_{1},S_{2},\ldots,S_{N}\},\quad S_{i}\cap S_{j}=\varnothing\;\;\forall\,i\neq j,\quad\bigcup_{i=1}^{N}S_{i}=\mathcal{C}.$
A corpus scale $n\in\{1,\ldots,N\}$ is realized by activating $n$ shards; we use the canonical prefix $\mathcal{C}^{(n)}=\bigcup_{i=1}^{n}S_{i}$, so that $|\mathcal{C}^{(n)}|\propto n$. For a query $q$, the retriever operates on $\mathcal{C}^{(n)}$ to retrieve top-$k$ documents, segment them into chunks, rerank, and pass the top-$m$ chunks to the generator.
@@ 3.2 Generator
We consider a family of generators $\{M_{x}\}$, where $M_{x}$ denotes a model with parameter size $x$ drawn from the same architecture family. Each generator takes as input a fixed template consisting of the query and retrieved chunks.
@@ 3.3 Trade-off Formalization
We adopt a full-factorial design pairing each corpus scale $n\in\{1,\ldots,N\}$ with each generator $M_{x}$. Throughout, we keep retrieval, prompting, and decoding settings fixed, so that only corpus scale $n$ and model size $x$ vary. Let $P_{m}(n,x)$ denote the evaluation score under metric $m\in\{\text{F1},\text{EM}\}$.
@@
$n^{\star}(x_{small}\!\to\!x_{large})\;:=\;\min_{m\in\{\text{F1},\,\text{EM}\}}\;\min\big\{n\,:\,P_{m}(n,x_{small})\,\geq\,P_{m}(1,x_{large})\big\},$
the smallest corpus scale where a smaller generator $M_{x_{small}}$ matches the 1-shard baseline ($n=1$) of a larger generator $M_{x_{large}}$. We report $n^{\star}$ and efficiency curves across $(n,x)$ to characterize corpus–generator trade-offs. Section 4 details datasets, metrics, and constants.
@@ 4.1 Benchmarks
Scope and limitation. Our evaluation targets open-domain QA with a large web corpus (ClueWeb22-A), which may overlap with modern LLM pretraining data; thus our results should be interpreted within this open-domain setting. Although our later analyses condition on questions that are initially incorrect without retrieval to isolate retrieval-driven gains, this setting still differs from domain-gap or enterprise RAG where the corpus is largely unseen during pretraining; thus, extrapolations should be made with caution.
@@ 4.2 Evaluation Metrics
We report Exact Match (EM) and token-level F1 using the official evaluation scripts. EM is the percentage of predictions that exactly match any normalized gold answer string. F1 is the average token-overlap F1 between the prediction and the gold answers, taking the maximum over gold answers for each question. Our analysis primarily focuses on these two metrics throughout the paper.
@@ 4.3 Retriever: Implementation Details
Retrieval pipeline. We use MiniCPM-Embedding-Light [10] for dense passage encoding and build ANN indices with DiskANN [26], which supports fast multi-shard retrieval and has been used in other large-scale ClueWeb22-A retrievers [5]. As an auxiliary check, we also encoded the corpus with a similarly sized Qwen3-Embedding-0.6B encoder [35] and observed consistent corpus-scaling trends. For each corpus scale $n$, the retriever searches the active shards to retrieve the top-10 documents, segments them into overlapping chunks, and reranks the candidates; we then pass the top-8 chunks to the generator. For reranking, we use the higher-capacity MiniCPM-Embedding model from the same embedding family to improve ranking quality, while keeping all retrieval and reranking settings fixed across $n$ to isolate corpus scaling effects.
@@ 4.4 Generator: Implementation Details
We instantiate $\{M_{x}\}$ using the open-source Qwen3 base series [33]: Qwen3-0.6B, 1.7B, 4B, 8B, and 14B. These models share the same tokenizer and a consistent architecture family as described in the Qwen3 technical report, spanning over an order of magnitude in parameter scale.
All models use identical prompting templates and decoding settings, enabling controlled comparisons of corpus–model trade-offs. While Qwen3 sizes share the same tokenizer and architecture family, different scales may involve minor training-recipe differences; thus, our results should be interpreted as empirical trends across sizes within the same Qwen3 series. We observed consistent trends in preliminary experiments with Qwen2.5 and early LLaMA models, but omit them due to the lack of a similarly homogeneous series and slower inference.
@@ 5.1 The Effect of Corpus Scaling as Compensation
We examine whether enlarging the retriever corpus can compensate for smaller LLMs, enabling them to match or surpass larger model. We parameterize corpus scale by the number of active shards $n$, where each shard indexes $\sim$22M documents from ClueWeb22-A; scaling the corpus corresponds to increasing $n$.
@@
$\times$8 34.72 24.19 42.12 31.09 46.91 34.60 48.04 35.73 49.34 36.80
$\times$9 34.49 23.97 42.04 30.75 46.64 34.31 48.27 35.61 49.36 36.63
$\times$10 35.35 24.59 42.86 31.32 47.08 34.43 48.62 36.07 49.75 37.37
$\times$11 35.48 25.10 42.78 30.98 47.43 34.71 49.06 36.01 50.38 37.59
$\times$12 35.56 25.10 43.14 31.32 47.89 34.88 48.73 35.78 50.43 37.70
We evaluate five Qwen3 variants ($M_{0.6B}$, $M_{1.7B}$, $M_{4B}$, $M_{8B}$, $M_{14B}$) and, for each model, vary $n$ by cumulatively activating shards under the same retrieval and prompting protocol, enabling controlled cross-model comparisons.
@@
Our results show clear evidence that corpus expansion enables smaller models to match or even outperform larger counterparts. On NQ, as shown in Table 1, we find that the smallest model needs more corpus to surpass the larger model $n^{\star}(0.6\text{B}\!\to\!1.7\text{B})=6$. For larger models, it is much easier: $n^{\star}(4\text{B}\!\to\!8\text{B})=2$ and $n^{\star}(8\text{B}\!\to\!14\text{B})=2$. These results indicate that, under our setup, scaling corpus size can be a more effective and efficient lever than simply scaling LLM size. Figures 2 and 2 visualize these catch-up points.
@@
$n^{\star}$ NQ TriviaQA WebQ NQ Qwen3-Emb
@@
[…]
5.1.2 Serving-Time Compute Proxy.
We provide a simple serving-time compute proxy by estimating FLOPs per query for retrieval and generation. Under a representative setting (top-K=10K{=}10, L=5K=50L{=}5K{=}50, dot-product retrieval with D=1024D{=}1024; generation with $S{\approx}2048$ prompt tokens and $T{\approx}64$ output tokens), a single-shard DiskANN retrieval costs $\sim$7.168 MFLOPs/query, while Qwen3-0.6B and Qwen3-1.7B generation cost $\sim$1.736 TFLOPs/query and $\sim$5.953 TFLOPs/query, respectively. This orders-of-magnitude gap is already sufficient to show that, in our setup, arithmetic compute is dominated by generation rather than retrieval. We focus on inference-time compute because our experiments vary serving-side design choices such as corpus scaling and generator size. This proxy does not account for one-time costs such as LLM pretraining or index/shard construction.
@@
Shards are balanced in size but not perfectly uniform, since randomization was applied locally rather than globally. Consequently, shard order may retain residual crawl/popularity gradients (e.g., earlier shards being slightly more likely to contain higher-traffic or higher-quality pages). To probe sensitivity to corpus quality, we reverse the shard order at retrieval time, replacing $S_{1},\ldots,S_{n}$} with $S_{N-n+1},\ldots,S_{N}$}, which delays the “head” content. As expected, this yields a modest drop in absolute end-task QA performance (Figure 3, left).
@@
In other words, lower average corpus quality shifts performance downward, yet the relative additional corpus needed for a smaller model to catch up with the next larger model remains largely stable. This supports our conclusion: corpus quality affects absolute accuracy, but scaling corpus quantity can still enable smaller generators to match or overtake larger ones with similar additional context. For consistency, we report all subsequent results using the forward order.
@@ 5.1.4 Why Does Corpus Scaling Improve RAG?
At the micro level, corpus scaling increases the likelihood that retrieved passages explicitly contain the gold answer. With a small corpus scale ($n=1$), retrieved chunks often lack factual mentions. The larger $n$ brings the direct answer terms into the context, providing the generator with grounding evidence as intended in RAG.
@@ 5.2.1 Classification Methodology
which captures the fraction of initially unanswerable questions solved at shard $n$. Note that although $\mathrm{CB}@n$ is computed cumulatively, it is an empirical statistic and need not be strictly monotonic in $n$ (e.g., additional shards may introduce noise).
@@
Figure 6 reveals a consistent and model-invariant pattern for how the scaling helps answer questions that are initially unanswerable without context, with similar trends observed on TriviaQA and WebQ (see the Appendix).
@@
5.2.2 Initial Jump, Subsequent Growth, and Saturation.
@@
Across all LLM sizes, adjacent-shard improvements in downstream QA (F1) exhibit a highly consistent pattern. The paired bootstrap differences from shard $n$ to n+1n{+}1 show large, reliably positive gains in the early steps, followed by rapid decay toward a near-zero regime (Figure 7). For all model sizes, the 95% CI typically begins to include zero around the $3{\to}4$ or $4{\to}5$ transition, indicating diminishing returns and that later fluctuations are often within statistical noise rather than systematic improvement; we observe the same conclusion when computing paired differences using EM.
@@
A consistent picture emerges for retrieval-driven gains on initially unanswerable questions. Corpus scaling yields a qualitatively similar CB trajectory across models: a sharp first jump, sustained gains up to roughly $n\!\approx\!6$, and diminishing returns thereafter. The per-shard increments $\Delta_{n}$ also follow a nearly identical pattern across models: peaking early and tapering to near zero (Figure 8). Together, these results suggest a size-invariant retrieval effect: additional shards do not yield systematically larger marginal gains for larger generators than for smaller ones. In practice, corpus expansion mainly shifts performance upward without materially changing the shape of the growth curve.
@@
As shown in Figure 9, the Utilization Ratio stays approximately constant across corpus scales and varies only slightly across models. Although both $\mathrm{CB}@n$ and $\mathrm{Coverage}@n$ increase with $n$, their ratio remains stable, indicating that corpus scaling primarily improves answer-bearing evidence coverage rather than changing how efficiently generators use the provided context. As a result, gains largely track the increased availability of relevant evidence as the corpus grows.